# Baikal-GVD Collaboration Measurement of the diffuse astrophysical neutrino flux over six seasons using cascade events from the Baikal-GVD expanding telescope

### v1

@@ Abstract
We present an updated measurement of the diffuse astrophysical neutrino flux using Baikal-GVD cascade data collected between April 2018 to March 2024. In this period, the detector grew from 15% to 55% of its baseline cubic kilometer configuration. The diffuse astrophysical neutrino flux is detected with a statistical significance of 5.1 . Assuming a single power law model of the astrophysical neutrino flux with identical contribution from each neutrino flavor, the following best-fit parameter values are found: the spectral index = 2.64 and the flux normalization = 4.42 per one flavor at 100 TeV. These results are broadly consistent with IceCube measurements.
@@ II BAIKAL-GVD NEUTRINO EXPERIMENT
Baikal-GVD is currently the largest neutrino telescope operating in the Northern Hemisphere [20]. The telescope is located in the Southern part of Lake Baikal ( N, E) at about 4 km from the shore. The detector is formed by sub-arrays, known as clusters, each of which is connected to the shore station by its own electro-optical cable. Each cluster is an independent array comprising 288 light sensors – optical modules (OMs), which are installed along vertical strings, with 36 OMs per string. The first full-scale Baikal-GVD cluster was deployed in April 2016. In 2017–2025 thirteen additional clusters were deployed and commissioned, increasing the total number of optical modules to over 4100 OMs (see Fig.1). The design and basic characteristics of the telescope data acquisition system are described elsewhere [20, 21].
@@ III DATA ANALYSIS
The search for high-energy astrophysical neutrinos using cascade-like events includes the selection and reconstruction of high-energy showers generated in neutrino interactions in the telescope detection volume. The cuts on quality variables are optimized on Monte Carlo simulations and tuned using the data sample accumulated in 2016–2017. In this analysis we used the event selection and shower reconstruction procedures similar to those used in our previous analysis [18]. We select events with a large multiplicity of triggered OMs, at three or more strings, and requiring that hits have charges p.e. and satisfy the causality condition [21].
The shower vertex coordinates are reconstructed by minimization of function using the time information from the selected hits [22]. The shower energy () and direction () are reconstructed by applying the maximum-likelihood method with the use of the reconstructed shower coordinates. Poorly reconstructed events are rejected by applying cuts on quality parameters, including the values of and maximum-likelihood function and OMs hit multiplicity [23]. The precision of the reconstruction of shower energy and direction was estimated using Monte Carlo simulation of the Baikal-GVD cluster [24]. The precision of energy reconstruction is typically varied in the range 10–20%. The precision of reconstruction of the shower direction is 2∘–4∘ (median value) [24]. We used Baikal-GVD Monte Carlo simulation packages to simulate the cosmic ray background with CORSIKA 7.74 [25] using the proton spectrum proposed in [26]. The primary interaction of protons in the Earth atmosphere was simulated based on the SIBYLL 2.3c model [27]. The propagation of muons in water down to the detector depth was simulated based on the MUM program [28]. The efficiency of neutrino event registration was estimated by simulating the passage of neutrinos through the Earth and the interaction in the sensitive volume of the facility using the neutrino cross sections from [29, 30], the lepton decay cross sections from [31], and the model of the Earth profile from [32]. The telescope response to the Cherenkov radiation of showers from neutrino interactions was simulated accounting for the shower development in water, as well as light absorption, scattering and light velocity dispersion in water.
Astrophysical neutrino event selection efficiencies were tested assuming a flux with equal numbers of neutrinos and antineutrinos, and with an equal neutrino flavour mixture at Earth: . The one flavour (1f) flux presented by Baikal-GVD in [18] was chosen as baseline:
The conventional atmospheric neutrino flux from pion and kaon decays was modeled according to [33]. Atmospheric prompt neutrino were simulated according to the BERSS model [34].
@@ IV RESULTS AND DISCUSSION
We use Baikal-GVD data collected between April 2018 and March 2024 for the search for astrophysical neutrinos. The telescope operated with 3 clusters during April 2018–March 2019, 5 clusters in April 2019–March 2020, 7 cluster in April 2020–March 2021, 8 clusters in April 2021–March 2022, and 10 clusters in April 2022–early 2023. From April 2023 to March 2024, the configuration included 11 complete clusters and one incomplete cluster. In this study, we report on results of data analysis for individual clusters as independent setups. A sample of 5.491010 events was collected by the basic trigger of the telescope. After applying noise hit suppression procedures, cascade reconstruction and cuts on reconstruction quality parameters, a sample of 12077 cascades with reconstructed energy TeV and OM hit multiplicity 11 was selected. It was shown in [18] that restricting the analysis to upward-going directions allows for effective suppression of the atmospheric muon background, thus improving the neutrino sample purity and enabling the extension of the analysis towards lower energies. Cascade-like events with reconstructed energy 15 TeV, OM hit multiplicity 11 and reconstructed zenith angle -0.25 were selected as astrophysical neutrino candidates. A total of 25 events have been selected in the April 2018–March 2024 data sample. The fraction of background events from atmospheric muons and neutrinos in the selected with this cuts sample is expected at a level of 20%. Finally, after applying additional cuts which suppress events from atmospheric muons and atmospheric muon neutrinos [35], 18 events have been selected, while 2.81 atmospheric background events are expected (1.9 from atmospheric conventional and prompt neutrinos and 0.9 events from mis-reconstructed atmospheric muons). The effect of the uncertainty of the detector response to signal and background is evaluated by varying input parameters in the Monte Carlo simulations. The uncertainty of the light absorption length is about 10%. Such variations lead to a change in the detection efficiency of high-energy cascades by about 18%-20% and shift the energy scale in the logarithm of the cascade energy by about 0.05. At the same time, the uncertainty of the reconstruction of the cascade direction weakly depends on the changes in the light absorption length. The optical module sensitivity varies within 10%. Also a 15% uncertainty on the normalization of the conventional atmospheric neutrino component is considered [33]. It is shown below that the total statistical and systematic uncertainty of the best fit astrophysical flux normalization is about 40%-50%. Given an upper limit on the prompt atmospheric neutrino flux normalization about 5.0, as it was assumed in IceCube’s 6 yr cascade analysis [15], an uncertainty of the prompt neutrino flux normalization would cause the uncertainty of the astrophysical best fit flux normalization less than 4%. For this reason, this uncertainty was not included in this analysis. The uncertainties coming from independent sources are added in quadrature in the overall estimation. Taking into account the systematic effects according to the method of [36], the significance of the excess was estimated to be 5.1 with the chance probability 2.1. The parameters of the 18 upward-going cascades are shown in Tab. I. The median value of the error in the reconstruction of the cascade direction varies from 1.6∘ to 5.4∘.
@@
where = 100 TeV. The best fit parameters for the observed data are determined by a binned likelihood approach. In this procedure, the data sample is binned in reconstructed shower energy. The observed count in each bin is compared to a model that predicts the mean count rate in each bin using a Poisson likelihood function:
The expected rates are composed by astrophysical neutrinos and background events of atmospheric muons and atmospheric neutrinos. The MC simulated templates of the cosmic signal and of the atmospheric backgrounds with different water parameters and OM efficiency were used for estimation. Accordingly, these templates include effects of change of detection efficiency and energy scale shift. Based on the Poisson likelihood function (3) the following test statistic (TS) is used to compare the observed event counts with the Monte Carlo model predictions:
The second term in (4) accounts for the systematic uncertainties discussed above which are incorporated in the test statistic as nuisance parameters in form of Gaussian distributions of prior and width deviation from central value . A maximum-likelihood method is applied to find the best-fit values of and by varying these parameters until (4) is minimized. We find the best-fit parameters as follows: the spectral index = 2.64 and the flux normalization for each neutrino flavour at = 100 TeV = 4.42.

### v2

@@ Abstract
We present an updated measurement of the diffuse astrophysical neutrino flux using Baikal-GVD cascade data collected between April 2018 to March 2024. In this period, the detector grew from 15% to 55% of its baseline cubic kilometer configuration. The diffuse astrophysical neutrino flux is detected with a statistical significance of 5.1 . Assuming a single power law model of the astrophysical neutrino flux with flavor equipartition, the following best-fit parameter values are found: the spectral index = 2.64 and the flux normalization = 4.42 per one flavor at 100 TeV. These results are broadly consistent with IceCube measurements.
@@ II BAIKAL-GVD NEUTRINO EXPERIMENT
Baikal-GVD is currently the largest neutrino telescope operating in the Northern Hemisphere [20], with a detection volume of about 0.7 km3 achieved in April 2025. The telescope is located in the Southern part of Lake Baikal ( N, E) at about 4 km from the shore and at the depth of deployment of 1366 m. The detector is formed by sub-arrays (clusters) of optical modules (OM) instrumented with 10-inch high-quantum-efficiency PMT HAMAMATSU36 R7081-100 and various sensors. Each cluster comprises 288 OM distributed on 8 vertical strings and hosted between depths of 750 and 1275 meters. The connection of each cluster to the shore station by its own electro-optical cable provides both independent detection of events by single cluster and by multi-clusters in their time-synchronized operations.
The first full-scale Baikal-GVD cluster was deployed in April 2016. After winter expedition in 2025 the telescope incorporates 14 clusters (see Fig.1) including in total 117 strings carrying 4212 OMs. The design and basic characteristics of the telescope data acquisition system are described elsewhere [20, 21].
@@ III DATA ANALYSIS
The search for high-energy astrophysical neutrinos is based on the selection and reconstruction of high-energy showers generated in the telescope detection volume. In this analysis we used the event selection and shower reconstruction procedures similar to those used in our previous analysis [18]. We select events with a multiplicity of triggered OMs, at three or more strings, and requiring that hits have charges p.e. and satisfy the causality condition [21].
The shower vertex coordinates are reconstructed by minimization of function using the time information from the selected hits [22]. The shower energy () and direction () are reconstructed by applying the maximum-likelihood method with the use of the reconstructed shower coordinates. Poorly reconstructed events are rejected by applying cuts on quality parameters, including the values of and maximum-likelihood function and OMs hit multiplicity [23]. The precision of energy reconstruction is typically varied in the range 10–20%. The precision of reconstruction of the shower direction is 2∘–4∘ (median value) [24]. The cosmic ray background was simulated with CORSIKA 7.74 [25] using the proton spectrum proposed in [26] and the SIBYLL 2.3c interaction model [27]. The propagation of muons in water was simulated based on the MUM program [28]. The passage of neutrinos through the Earth and the interaction in the sensitive volume of the telescope were simulated using the neutrino cross sections from [29, 30], the lepton decay cross sections from [31], and the model of the Earth profile from [32]. The telescope response to the Cherenkov radiation of showers from neutrino interactions was simulated accounting for the shower development in water, as well as light absorption, scattering and light velocity dispersion in water.
@@ IV RESULTS AND DISCUSSION
We use Baikal-GVD data collected between April 2018 and March 2024 for the search for astrophysical neutrinos. A season of data taken covers temporal period from April to next March. The telescope operated with 3 clusters in 2018–2019 season, 5 clusters in 2019– 2020 season, 7 in 2020–2021 season, 8 clusters in 2021–2022, and 10 clusters in April 2022–early 2023. From April 2023 to March 2024, the configuration included 11 complete clusters and one incomplete cluster. In this study, we report on results of data analysis for individual clusters as independent setups. A sample of 5.491010 events was collected by the basic trigger of the telescope. After applying noise hit suppression procedures, cascade reconstruction and cuts on reconstruction quality parameters, a sample of 12077 cascades with reconstructed energy TeV and OM hit multiplicity 11 was selected. It was shown in [18] that restricting the analysis to upward-going directions allows for effective suppression of the atmospheric muon background, thus improving the neutrino sample purity and enabling the extension of the analysis towards lower energies. Cascade-like events with reconstructed energy 15 TeV, OM hit multiplicity 11 and reconstructed zenith angle -0.25 were selected as astrophysical neutrino candidates. A total of 25 events have been selected in the April 2018–March 2024 data sample. The fraction of background events associated with atmospheric muons and neutrinos in the sample selected using these cuts is expected at a level of 20%. Finally, after applying additional cuts which suppress events from atmospheric muons and atmospheric muon neutrinos [35], 18 events have been selected, while 2.81 atmospheric background events are expected (1.9 from atmospheric conventional and prompt neutrinos and 0.9 events from mis-reconstructed atmospheric muons). The effect of the uncertainty of the detector response to signal and background is evaluated by varying input parameters in the Monte Carlo simulations. The uncertainty of the light absorption length is about 10%. Such variations lead to a change in the detection efficiency of high-energy cascades by about 18%-20% and shift the energy scale in the logarithm of the cascade energy by about 0.05. At the same time, the uncertainty of the reconstruction of the cascade direction weakly depends on the changes in the light absorption length. The optical module sensitivity varies within 10%. Also a 15% uncertainty on the normalization of the conventional atmospheric neutrino component is considered [33]. It is shown below that the total statistical and systematic uncertainty of the best fit astrophysical flux normalization is about 30%-50%. Given an upper limit on the prompt atmospheric neutrino flux normalization about 5.0, as it was assumed in IceCube’s 6 yr cascade analysis [15], an uncertainty of the prompt neutrino flux normalization would cause the uncertainty of the astrophysical best fit flux normalization less than 4%. For this reason, this uncertainty was not included in this analysis. The uncertainties coming from independent sources are added in quadrature in the overall estimation. Taking into account the systematic effects according to the method of [36], the significance of the excess was estimated to be 5.1 with the chance probability 2.1. The parameters of the 18 upward-going cascades are shown in Tab. I. The median value of the error in the reconstruction of the cascade direction varies from 1.6∘ to 5.4∘.
@@
where = 100 TeV. By the same way as in our previous analysis [18] the best fit parameters for the observed data are determined by a binned likelihood approach. In this procedure, the data sample is binned in reconstructed shower energy. The observed count in each bin is compared to a model that predicts the mean count rate in each bin using a Poisson likelihood function. The expected rates are composed by astrophysical neutrinos and background events of atmospheric muons and atmospheric neutrinos. The MC simulated templates of the cosmic signal and of the atmospheric backgrounds with different water parameters and OM efficiency were used for estimation. Accordingly, these templates include effects of change of detection efficiency and energy scale shift. Following the Poisson likelihood function (3) and the compiled test statistic (TS) as in our previous study [18], we fitted the observed event counts by the Monte Carlo model predictions. The systematic uncertainties discussed above were incorporated in the test statistic as () nuisance parameters in form of Gaussian distributions of prior and width deviation from central value . A maximum-likelihood method is applied to find the best-fit values of and by varying these parameters until TS is minimized. We find the best-fit parameters as follows: the spectral index = 2.64 and the flux normalization for each neutrino flavour at = 100 TeV = 4.42.